OpenStudy (anonymous):

Find the value for the discriminant for the following equation: (2x +1)(x – 3) = 0

zepdrix (zepdrix):

Hey Jacky :) Do you understand how to multiply out the bracketed terms?

zepdrix (zepdrix):

We need this in standard form before we can look at the discriminant: \(\Large\rm ax^2+bx+c=0\)

OpenStudy (anonymous):

Yes. I do. I would get 2x^2-5x-3

zepdrix (zepdrix):

Mmm ok great!

OpenStudy (anonymous):

I am confused on what I should do next.

zepdrix (zepdrix):

Our quadratic formula allows us to solve for x, \[\Large\rm x=\frac{-b\pm\sqrt{\color{orangered}{b^2-4ac}}}{2a}\] The discriminant is the orange part, that's all we're concerned with here.

OpenStudy (anonymous):

I got 7, is that correct?

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{b^2-4ac}\]Usually we want to know if this `greater than` `equal to` or `less than` zero, to find out information about the polynomials roots.

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{b^2-4ac}=(-5)^2-4(2)(-3)\]

zepdrix (zepdrix):

Hmm 7 doesn't look quite right :o

OpenStudy (anonymous):

so would it be 49? I do not square root the b^2-4ac part?

zepdrix (zepdrix):

Correct :) No square root needed. They just want the `value` of the discriminant (stuff under the root). Yay good job \c:/

OpenStudy (anonymous):

Alright, thanks! I understand now. :)

zepdrix (zepdrix):

cool